# Corks, exotic 4-manifolds and genus functions

### v3

@@ 11.2. Properties
We show that the stabilities of algebraic inequivalences also hold for $G$ invariants. More precisely, we prove the following theorem for attachments of (quasi-)invertible cobordisms.
@@
$(1)$ Theorems 7.1, 7.8, 7.9 and 7.10 and Corollaries 7.4, 7.5 and 7.7 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $P$ is invertible.
$(2)$ Suppose that $G$ is preserved under attachments of $3$- and $4$-handles. Then, Theorems 7.1, 7.8, 7.9 and 7.10 and Corollaries 7.4, 7.5 and 7.7 hold even when “genus functions” in each statement is replaced with “$G$ invariants”.
@@
Due to Theorem 11.10.(2), we see that the left inequality in Proposition 7.6 holds for $G$ invariants by the identical argument. Therefore, we can prove the claims of (2) by the identical arguments. Due to Theorem 11.10.(1), we can prove the claims of (1) in the same way. ∎
@@
$(1)$ For boundary sums, Theorems 8.4, 8.8, 8.9, 8.10 and 8.11 and Corollary 8.6 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $L$, $L_{1}$ and $L_{2}$ are empty sets.
$(2)$ Suppose that $G$ is preserved under attachment of a $4$-handle. Then, Theorems 8.4, 8.8, 8.9, 8.10 and 8.11 and Corollary 8.6 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $L$, $L_{1}$ and $L_{2}$ are empty sets. If furthermore $G$ is preserved by attachment of a $3$-handle, then these claims and Proposition 8.13 for $G$ invariants hold without any condition on $L$, $L_{1}$ and $L_{2}$.
@@
Due to Theorem 11.10, we see that the second claim of Theorem 8.4 and the left inequality in the second claim of Proposition 8.5 hold for $G$ invariants by the identical arguments. Therefore, we can prove the claim (1) by the identical arguments. In the case where $G$ is preserved under attachment of a $4$-handle, we can prove the claim (2) in the same way. We remark that the left inequality in the first claim of Proposition 8.5 also holds for $G$ invariants in this case. ∎
The above theorems imply that if algebraic equivalent classes of $G$ invariants can detect a family of exotic 4-manifolds, then they can detect infinitely many families of exotic 4-manifolds.
@@
$(1)$ If the members $X_{i}$’s have non-empty boundary, then there exists a doubly indexed family $\{X_{i,j}\}_{(i,j)\in A\times\mathbb{Z}}$ of $4$-manifolds having the following properties.
@@
$(2)$ If $G$ is preserved by a $4$-handle, then the conclusion of the above $(1)$ holds even in the case where $X_{i}$’s are closed $4$-manifolds.
@@
We first consider the case (1). By Theorems 11.12 and 11.13, we can construct many examples of desired families, for example, by taking boundary sums of $X_{i}$’s with $S^{1}\times D^{3}$’s. Their homeomorphism types are detected by their first homology groups. In the (2) case, we can similarly construct many examples of desired families, for example, by taking connected sums of $X_{i}$’s with $S^{1}\times D^{3}$’s. We remark that, if $G$ is furthermore preserved by a 3-handle, then we can give desired families of closed 4-manifolds as well, by taking connected sums with $S^{4}(L)$’s. ∎
In contrast to the above corollary, we also show that there exist various infinite families of exotic 4-manifolds which can not be detected by their $G$ invariants. More precisely, we show the following theorem.
@@
Proposition 9.3 and Theorems 9.4 and 9.5 hold even when “genus functions” in each statement is replaced with “$G$ invariants”.
Proof.
Due to Theorem 11.10, the proofs are identical with the original ones. Note that the boundary of a 4-manifold is assumed to be connected as in the original claims. ∎
@@
$(1)$ The topological genus function of a 4-manifold is defined similarly to the (smooth) genus function by using locally flat embedded surfaces instead of smoothly embedded surfaces, and this function also has long been studied, see [57] and the references therein, for example. Topological genus functions clearly form an (unoriented) diffeomorphism invariant of genus function type. Due to the transversality theorem for locally flat proper submanifolds (see [15, Theorem 1.5] and [57]), the proof of Lemma 7.2 shows that topological genus functions are preserved by 3- and 4-handles. Since two homeomorphic 4-manifolds have equivalent topological genus functions, they cannot be used for detecting exotic 4-manifolds. However, results in this section for topological genus functions might be useful for determining their values and also for detecting homeomorphism types by using stabilities of algebraic inequivalences.
$(2)$ Due to our convention, an invariant $G$ of genus function type is defined only for compact oriented 4-manifolds, but $G$ can be defined for non-compact 4-manifolds as well at least for genus functions. For such an invariant $G$, it easily follows from the definition of genus function type that the $G$ invariants of a 4-manifold and its interior are algebraically equivalent. Thus, for any two 4-manifolds having algebraically inequivalent $G$ invariants, their interiors are not orientedly diffeomorphic.

### v5

@@ 11.2. Properties
Using oriented diffeomorphism invariants of genus function type, we can characterize $S^{4}$ and its submanifolds as follows.
Proposition 11.10.
$(1)$ A $4$-manifold is diffeomorphic to $S^{4}$ if and only if any oriented diffeomorphism invariant of genus function type is preserved under connected sum with the $4$-manifold.
$(2)$ A $4$-manifold with non-empty boundary admits an embedding into $S^{4}$ if and only if any oriented diffeomorphism invariant of genus function type is preserved under boundary sum with the $4$-manifold.
Proof.
The “only if” part of the claim (1) is straightforward, and that of (2) is Theorem 11.8. In the rest, we prove the “if” part of each claim.
(1) For a 4-manifold $X$ and a class $\alpha\in H_{2}(X)$, we set $D^{S^{4}}_{X}(\alpha)=0$ (resp. $D^{S^{4}}_{X}(\alpha)=\infty$) in the case where $X$ is diffeomorphic (resp. not diffeomorphic) to $S^{4}$. Then, the resulting map $D^{S^{4}}_{X}:H_{2}(X)\to\{0,\infty\}$ gives an oriented diffeomorphism invariant of genus function type. Now let $Z$ be a 4-manifold such that any oriented diffeomorphism invariant of genus function type is preserved under connected sum with $Z$. Then, $Z$ satisfies $D^{S^{4}}_{S^{4}\#Z}(0)=D^{S^{4}}_{S^{4}}(0)=0$, and hence $Z$ is diffeomorphic to $S^{4}$. Therefore, the claim (1) follows.
(2) For a 4-manifold $X$ and a class $\alpha\in H_{2}(X)$, we set $E^{S^{4}}_{X}(\alpha)=0$ (resp. $E^{S^{4}}_{X}(\alpha)=$-$\infty$) in the case where $X$ admits an embedding (resp. no embedding) into $S^{4}$ that preserves the orientations. Then, the resulting map $E^{S^{4}}_{X}:H_{2}(X)\to\{-\infty,0\}$ gives an oriented diffeomorphism invariant of genus function type. Now let $Z$ be a 4-manifold with non-empty boundary such that any oriented diffeomorphism invariant of genus function type is preserved under boundary sum with $Z$. Then, $Z$ satisfies $E^{S^{4}}_{D^{4}\natural Z}(0)=E^{S^{4}}_{D^{4}}(0)=0$, and hence $Z$ admits an embedding into $S^{4}$. Therefore, the claim (2) follows. ∎
Remark 11.11.
(1) Similarly to the above proof, we see that for any closed 4-manifold, its diffeomorphism type and oriented diffeomorphism type are determined by oriented diffeomorphism invariants of genus function type similar to $D^{S^{4}}$.
(2) For a closed (possibly disconnected) 3-manifold $M$, we define an oriented diffeomorphism invariant $E^{I\times M}$ in the same way as the above $E^{S^{4}}$. Using this invariant similarly to the above proof, we can prove that a cobordism with negative boundary $M$ admits an embedding into $I\times M$ that preserves the orientations if any oriented diffeomorphism invariant of genus function type is preserved under attachment of the cobordism.
We show that the stabilities of algebraic inequivalences also hold for $G$ invariants and their torsion-free versions. More precisely, we prove the following theorem for attachments of (quasi-)invertible cobordisms.
@@
$(1)$ Theorems 7.9, 7.10 and 7.11 and Corollaries 7.4, 7.5 and 7.6 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where the cobordism $P$ in Subsection 7.1 is invertible.
$(2)$ Suppose that $G$ is preserved under attachment of a $3$- and a $4$-handle. Then, Theorems 7.9, 7.10 and 7.11 and Corollaries 7.4 and 7.6 hold even when “genus functions” in each statement is replaced with “$G$ invariants”.
@@
Due to Theorem 11.8, we see that the left inequality in Theorem 7.7 and Corollary 7.8 holds for $G$ invariants under the above assumptions by the identical argument. Therefore, we can prove the claims by the identical arguments. ∎
@@
$(1)$ For boundary sums, Theorems 8.7, 8.8, 8.9 and 8.10 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $L$, $L_{1}$ and $L_{2}$ are empty sets.
$(2)$ Suppose that $G$ is preserved under attachment of a $4$-handle. Then, Theorems 8.7, 8.8, 8.9 and 8.10 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $L$, $L_{1}$ and $L_{2}$ are empty sets. If furthermore $G$ is preserved under attachment of a $3$-handle, then these claims and Theorem 8.12 for $G$ invariants hold without any condition on $L$, $L_{1}$ and $L_{2}$.
@@
Due to Theorem 11.8, we see that the left inequality in the second claim of Theorem 8.5 and the second claim of Corollary 8.6 hold for $G$ invariants by the identical arguments. Therefore, we can prove the claim (1) by the identical arguments. We can similarly prove the claim (2) under the above assumptions. ∎
The above theorems imply that if algebraic equivalent classes of $G$ invariants can detect a family of orientedly exotic 4-manifolds, then they can detect infinitely many families of orientedly exotic 4-manifolds.
@@
$(1)$ If the members $X_{i}$ have non-empty boundary, then there exists a doubly indexed family $\{X_{i,j}\}_{(i,j)\in A\times\mathbb{Z}}$ of $4$-manifolds having the following properties.
@@
$(2)$ If $G$ is preserved under attachment of a $4$-handle, then the conclusion of the above $(1)$ holds even in the case where each $X_{i}$ has no boundary.
$(3)$ If $G$ is preserved under attachment of a $3$- and a $4$-handle, and each $X_{i}$ has no boundary, then the conclusion of the above $(1)$ still holds, and furthermore the members $X_{i,j}$ can be chosen to be closed $4$-manifolds.
@@
We first consider the case (1). By Theorems 11.12 and 11.13, we can construct many examples of desired families, for example, by taking boundary sums of the 4-manifolds $X_{i}$ with codimension-0 submanifolds of $S^{4}$. See also Lemma 8.1. In the (2) case, we can similarly construct many examples of desired families by taking connected sums instead of boundary sums. In the (3) case, we can give desired families of closed 4-manifolds by taking connected sums with $S^{4}(L)$ for an $S^{2}$-link $L$ in $S^{4}$. ∎
As shown in Remark 11.11, every exotic pair of closed 4-manifolds admits an oriented diffeomorphism invariant of genus function type that can detect the pair. By contrast, we show that there exist various infinite families of exotic 4-manifolds which can not be detected by any such invariant.
@@
For any $4$-dimensional 22-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ such that, for any oriented diffeomorphism invariant $G$ of genus function type, their $G$ invariants are pairwise equivalent.
Theorem 11.16.
Suppose that a $4$-manifold $X$ with connected boundary admits an embedding into a symplectic $4$-manifold with weakly convex boundary. Then, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ such that, for any oriented diffeomorphism invariant $G$ of genus function type, their $G$ invariants are pairwise equivalent and are algebraically equivalent to that of $X$.
Proof of Theorems 11.15 and 11.16.
Due to Theorem 11.8, the proofs are identical with those of Theorems 9.4 and 9.5. ∎
We close this section by making a few remarks.
@@
$(1)$ The topological genus function of a 4-manifold is defined similarly to the (smooth) genus function by using locally flat embedded surfaces instead of smoothly embedded surfaces, and this function also has long been studied, see [59] and the references therein, for example. Topological genus functions clearly form an (unoriented) diffeomorphism invariant of genus function type. Due to the transversality theorem for locally flat proper submanifolds (see [16, Theorem 1.5] and [59]), the proof of Lemma 7.2 shows that topological genus functions are preserved under attachment of a 3- and a 4-handle. Since two homeomorphic 4-manifolds have equivalent topological genus functions, they cannot be used for detecting exotic 4-manifolds. However, results in this section for topological genus functions might be useful for determining their values and also for detecting homeomorphism types by using stabilities of algebraic inequivalences.
$(2)$ Due to our convention, we defined the notion of genus function type only for (oriented) diffeomorphism invariants of compact 4-manifolds, but in the same manner, we can define the notion more generally for invariants of 4-manifolds which are not necessarily compact. For example, the genus function and the invariants $E^{SW}$ and $g^{\mathrm{dp}}$ defined in Example 11.5 are of genus function type in the latter sense. For such an invariant $G$, we can show that the $G$ invariants of a 4-manifold and its interior are algebraically equivalent similarly to the proof of Theorem 7.1. Thus, for any two 4-manifolds having algebraically inequivalent $G$ invariants, their interiors are not orientedly diffeomorphic. On the other hand, we can also define the notion of genus function type for (oriented) diffeomorphism invariants of compact 4-manifolds having connected (possibly empty) boundary. For such invariants, our results in this section still hold for 4-manifolds with connected (possibly empty) boundary, as easily seen from the proofs.